### v1

# Multi-Agent Guided Policy Search for Non-Cooperative Dynamic Games

Abstract
Multi-agent reinforcement learning (MARL) optimizes strategic interactions in non-cooperative dynamic games, where agents have misaligned objectives. However, data-driven methods such as multi-agent policy gradients (MA-PG) often suffer from instability and limit-cycle behaviors. Prior stabilization techniques typically rely on entropy-based exploration, which slows learning and increases variance. We propose a model-based approach that incorporates approximate priors into the reward function as regularization. In linear quadratic (LQ) games, we prove that such priors stabilize policy gradients and guarantee local exponential convergence to an approximate Nash equilibrium. We then extend this idea to infinite-horizon nonlinear games by introducing Multi-agent Guided Policy Search (MA-GPS), which constructs short-horizon local LQ approximations from trajectories of current policies to guide training. Experiments on nonlinear vehicle platooning and a six-player strategic basketball formation show that MA-GPS achieves faster convergence and more stable learning than existing MARL methods.
I Introduction
Non-cooperative dynamic games [1] provide a principled framework for modeling strategic interactions among self-interested agents, with applications in robotic coordination [2], autonomous driving [3], and human–robot interaction [4]. A central solution concept is the Nash equilibrium [1], which captures unilateral optimal decision-making among agents. Multi-agent reinforcement learning (MARL) has emerged as a promising data-driven tool for computing Nash equilibria in high-dimensional games. However, learning Nash equilibria remains challenging due to non-stationarity from simultaneous policy updates and conflicting objectives of agents. As a result, multi-agent policy gradients (MA-PG) often exhibit unstable dynamics or converge to limit cycles [5], even under smooth objectives and dynamics.
Classical approaches instead rely on model-based optimization [1, 6]. For example, [7] and [8] compute local Nash equilibrium using iterative linear quadratic (LQ) game approximations and Newton-type algorithms. While these methods usually converge stably, they scale poorly as the number of agents increases and the dynamics become nonlinear with complex objectives. In such cases, they become computationally expensive and may converge to inferior policies [9].
These limitations highlight the complementary strengths of MARL and mode-based optimization, suggesting that a hybrid method could provide both scalability and stability. In this paper, we move beyond prior efforts that integrate such approaches for cooperative multi-agent control [10, 11] and ask: how can we integrate these two approaches to improve policy learning for non-cooperative dynamic games?
Our key idea is to augment MA-PG methods with a model-based prior, defined as an approximate Nash equilibrium solution derived from local models (e.g., LQ approximations). This prior provides a shared reference that stabilizes policy updates across agents, avoids the limit cycles of pure MA-PG methods caused by non-unique Nash equilibria, simultaneous learning, and conflicting objectives, and ultimately guides agents toward a consistent Nash equilibrium.
We propose a framework that incorporates this prior into the reward function to regularize MA-PG updates. In doing so, it offers a new mechanism for stabilizing MARL that outperforms existing techniques based on entropy-driven exploration [12, 13]. Our method not only achieves faster convergence but also yields neural network policies that can be executed in real time for complex multi-agent dynamic games, where classical model-based optimization methods are computationally demanding for timely decision-making. Our contributions are:
  1. 1.
    For LQ games, characterized by linear dynamics and quadratic objectives, we theoretically prove that any stabilizing feedback policy, not necessarily to be a Nash equilibrium policy, can serve as effective guidance and establish formal guarantees on the stability and local convergence of guided policy gradients toward an approximate Nash equilibrium.
  2. 2.
    We extend the framework to nonlinear games by leveraging the fact that Nash equilibria of local LQ game approximations closely resemble those of the original nonlinear games. To construct these approximations efficiently, we use trajectories generated by the current neural network policies and treat the Nash equilibrium of the corresponding local LQ game approximation as guidance, eliminating the need for precomputed guidance policies used in classical guided policy search.
  3. 3.
    We evaluate our method on LQ games, nonlinear vehicle platooning, and a six-player strategic basketball formation, demonstrating faster and more stable convergence with higher-quality policies than state-of-the-art MARL methods in dynamic games with continuous state and action spaces.
II Related works
Model-based solutions to non-cooperative games. Computing Nash equilibria of non-cooperative dynamic games is NP-hard [14]; however, under smoothness assumptions on dynamics and costs, they can be solved in polynomial time [1]. For instance, LQ games enable efficient computation, with feedback Nash equilibria obtained via coupled Riccati equations or KKT conditions [8]. Extending these solutions to general nonlinear games remains difficult. Recent advances include iLQGames, which applies local LQ models and dynamic programming in an iterative scheme [7], and differential dynamic programming [6]. Other approaches compute equilibria using Newton’s method for KKT systems [15, 8] or approximate dynamic programming [16, 17], but high-dimensional nonlinear games remain challenging.
Data-driven MARL methods. MARL has been extensively studied in cooperative settings [18, 19, 20, 21, 11], with notable successes in actor–critic methods [22, 23, 24, 25], value decomposition [26, 27], and graph neural network approaches [28]. In contrast, learning game-theoretic equilibria [29] (e.g., Nash equilibria) with deep RL remains challenging due to coupled dynamics and conflicting objectives. Policy gradient methods converge reliably only in special classes such as two-agent zero-sum games [30, 31] and potential games [32], where agents’ costs are symmetric. In general non-cooperative games, however, they may converge to limit cycles rather than a Nash equilibrium [5]. Furthermore, directly applying guided policy search [33], which improves training in single-agent RL, to multi-agent Nash equilibrium settings is nontrivial due to misaligned objectives, except in fully cooperative cases [10, 11].
Training stability and variance reduction in deep RL. Introducing random noise into the dynamics can stabilize MA-PG methods in non-cooperative games [12, 13], suggesting that stochastic policies may improve MARL performance. However, this stabilization comes with increased gradient variance. Variance reduction techniques are well-established in single-agent RL [34], focusing on managing the bias-variance trade-off [35, 36, 37]. In multi-agent contexts, existing variance reduction methods include optimal baselines [38] and spectral normalization of critics [39], though these approaches do not incorporate model-based priors.
Model-based priors have been explored to reduce variance in RL [40]. Unlike [40], which regularizes the policy output directly, our method incorporates the regularization into the reward function. Prior works such as [12, 13, 41] stabilize MA-PG in finite-horizon games using entropy-driven exploration. In contrast, we show that suitable policy guidance ensures stability in infinite-horizon settings, which could be more practical for real-time decision-making since the optimal policy is time-invariant and avoids computing complex time-varying Nash equilibrium policies.
III Background: non-cooperative dynamic games
III-A Problem settings
We consider an N{N}-agent discrete-time non-cooperative dynamic game. At each stage t{t}, we associate with each agent $i\in[{N}]:=\{1,\dots,{N}\}$ a state vector ${x}_{t}^{i}\in\mathbb{R}^{{n}_{i}}$ and a control input vector ${u}_{t}^{i}\in\mathbb{R}^{{m}_{i}}$. Let ${n}:=\sum_{i=1}^{N}{n}_{i}$ be the dimension of the joint state, and ${m}:=\sum_{i=1}^{N}{m}_{i}$ be the dimension of the joint control input. We denote by ${x}_{t}:=[{x}_{t}^{i}]_{i=1}^{N}$ and ${u}_{t}=[{u}_{t}^{i}]_{i=1}^{N}$. Let $\mathcal{D}$} be the initial state distribution. Following standard assumptions in smooth dynamic games involving continuous state and action spaces [1, 7, 8], we consider the differentiable dynamics $f$:
${x}_{{t}+1}={f}({x}_{t},{u}_{t}^{1},\dots,{u}_{t}^{N}),\ \ {x}_{0}\sim{\mathcal{D}},\ \ t=0,1,\dots$ (1)
and each agent $i$ has an individual differentiable stage cost [^1] This assumption does not restrict the applicability of our method to games with non-differentiable objectives or dynamics; for instance, in the six-player strategic basketball formation experiments of Section VI, our method still outperforms existing MA-PG approaches despite the non-differentiable objective. ${c}^{i}({x}_{t},{u}_{t}^{1},\dots,{u}_{t}^{N})$. Let ${\pi}^{i}(x)$ be the time-invariant policy of agent $i$, which in general is a random distribution conditioned on the state variable $x$. Define ${\pi}({x}):=[{\pi}^{1}({x}),\dots,{\pi}^{N}({x})]$, and ${u}\sim{\pi}({x})$, where ${u}^{i}\in{\pi}^{i}({x})$, $\forall i\in[N]$. Denote by ${\pi}^{-i}:={\pi}\setminus\{{\pi}^{i}\}$. Given the policies $\{{\pi}^{i}\}_{i=1}^{N}$, we define the value function of the $i$-th agent ${V}^{i,{\pi}}(\cdot):\mathbb{R}^{n}\to\mathbb{R}$ as
${V}^{i,{\pi}}(x)=\mathbb{E}_{\begin{subarray}{c}{u}_{t}\sim{\pi}({x}_{t})\end{subarray}}\Big[\sum_{{t}=0}^{\infty}\gamma^{t}{c}^{i}({x}_{t},{u}^{1}_{t},\cdots{u}^{N}_{t})\Big],$ (2)
where $\gamma\in(0,1]$ is the time-discount factor. We have the Bellman equation for the $i$-th agent,
${V}^{i,{\pi}}({x})=\mathbb{E}_{\begin{subarray}{c}{u}_{t}\sim{\pi}({x}_{t})\end{subarray}}\big[{c}^{i}({x},{u}^{1},\cdots,{u}^{N})+\gamma{V}^{i,{\pi}}({f}({x},{u}^{1},\dots,{u}^{N}))\big].$
In a non-cooperative game, each agent $i\in{N}$ minimizes its cumulative cost ${J}^{i}$ over its policy ${\pi}^{i}$, while fixing ${\pi}^{-i}$
${J}^{i}({\pi}^{i},{\pi}^{-i}):=\mathbb{E}_{{x}_{0}\sim{\mathcal{D}}}\big[{V}^{i,{\pi}}({x}_{0})\big].$ (4)
A set of policies $\{{\pi^{i*}}\}_{i=1}^{N}$ is a local feedback Nash equilibrium if there exists $\epsilon>0$ such that for all $i\in[N]$, ${J}^{i}({\pi^{i}},{\pi}^{-i*})\geq{J}^{i}({\pi^{i*}},{\pi}^{-i*}),\ \forall{\pi^{i}},\ \text{s.t.}\;D_{KL}({\pi^{i}}|{\pi^{i*}})\leq\epsilon,$ where $D_{KL}$ is the KL divergence. For deterministic policies, this can be replaced with the L2 norm.
In MA-PG methods, each agent $i$’s policy is updated as:
${\pi}^{i}\leftarrow{\pi}^{i}-{\eta}\nabla_{{\pi}^{i}}{J}^{i}({\pi}^{i},{\pi}^{-i}),$ (5)
where $\eta$} is the step size. We denote the concatenation of all agents’ policy gradient as the pseudo-gradient $w({\pi})$:
$w({\pi}):=\begin{bmatrix}\nabla_{{\pi}^{1}}{J}^{1}({\pi}^{1},{\pi}^{-1})\\ \vdots\\ \nabla_{{\pi}^{{N}}}{J}^{N}({\pi}^{{N}},{\pi}^{-{N}})\end{bmatrix}$. (6)
We can represent the MA-PG update as a dynamical system with respect to $\pi$}, often referred to as the MA-PG dynamics:
${\pi}\leftarrow{\pi}-\eta\nabla w({\pi}).$ (7)
III-B Non-cooperative LQ games
In this subsection, we introduce a special class of non-cooperative games, LQ games [1], in order to present a relatively simple case where we can theoretically characterize the performance of policy gradient methods. In infinite-horizon LQ games, we consider
$\min_{{\pi}^{i}}\$$\mathbb{E}_{x_{0}\sim{\mathcal{D}}}\Big[\sum_{{t}=0}^{\infty}{x}_{t}^{\top}{Q}^{i}{x}_{t}+{u}_{t}^{\top}{R}^{i}{u}_{t}^{i}\Big]$ (8)
s.t. ${x}_{{t}+1}={A}{x}_{t}+{B}{u}_{t},\;{u}_{t}\sim{\pi}^{i}({x}_{t}),\;\forall t=0,1,\dots\vskip-2.5pt$
where discount factor is set to 1. Note that for policies resulting in stable closed-loop dynamics, we do not need to introduce the discount. More formally, we consider the following assumption:
Assumption 1.
The matrices $\{{Q}^{i},{R}^{i}\}_{i=1}^{N}$ are positive definite. The linear dynamics in (8) is stabilizable, i.e., there exists a set of time-invariant feedback matrices $\{{K}^{i}\}_{i=1}^{N}$ such that the closed-loop dynamics ${x}_{{t}+1}=\bar{A}_{K}{x}_{t}$ is exponentially stable, where $\bar{A}_{K}:={A}-\sum_{i=1}^{N}{B}^{i}{K}^{i}$. To ensure sufficient exploration, the covariance matrix $\Sigma_{0}:=\mathbb{E}_{x_{0}}[x_{0}x_{0}^{\top}]$ of the initial state distribution $\mathcal{D}$} is assumed to be full rank.
Assumption 1 is used only for the theoretical analysis of our method and is not required in practical experiments. It ensures the existence of a Nash equilibrium of (8). In LQ games with deterministic dynamics, we can specify the agent’s policy as a deterministic linear feedback policy, each given by $u_{t}^{i}=-K^{i}x_{t}$. With a slight abuse of notation $w(\cdot)$, we consider the following pseudo-gradient [5]
$w(K):=\begin{bmatrix}\nabla_{K^{1}}{J}^{1}(K^{1},\dots,K^{N})\\ \vdots\\ \nabla_{K^{N}}{J}^{N}(K^{1},\dots,K^{N})\end{bmatrix},K:=[K^{i}]_{i=1}^{N},$ (9)
to analyze the policy gradient method for computing the feedback Nash equilibrium. We analytically derive the policy gradient by substituting the linear policy into the Bellman equation (3),
${J}^{i}($ $K^{1},\dots,K^{N})=$ (10)
$\mathbb{E}_{x_{0}\sim{\mathcal{D}}}\Big[x_{0}^{\top}Q^{i}x_{0}+x_{0}^{\top}{K^{i}}^{\top}R^{i}K^{i}x_{0}+x_{0}^{\top}\bar{A}_{K}^{\top}P_{K}^{i}\bar{A}_{K}x_{0}\Big],$
where $P_{K}^{i}$ is given as the unique positive definite solution of the linear quadratic Bellman equation
$P_{K}^{i}=\bar{A}_{K}^{\top}P_{K}^{i}\bar{A}_{K}+{K^{i}}^{\top}R^{i}K^{i}+Q^{i}.\vskip-2.5pt$ (11)
Under Assumption 1, we have $\mathbb{E}_{x_{0}}\big[x_{0}x_{0}^{\top}\big]\succ 0$ and $\Sigma_{K}:=\mathbb{E}_{x_{0}}\big[\sum_{t=0}^{\infty}x_{t}x_{t}^{\top}\big]\succ 0$. As noted in [5], by taking the gradient of (10) with respect to $K^{i}$, we have
$\nabla_{K^{i}}{J}^{i}(K^{1},\dots,K^{N})=2(R^{i}K^{i}-{B^{i}}^{\top}P_{K}^{i}\bar{A}_{K})\Sigma_{K}.$ (12)
IV Analysis: Stability and bias of the guided policy optimization in LQ games
In this section, we present our analytical results for guided policy optimization in LQ games. First, we introduce a model-based guidance term into the original cost function, which serves as a regularization mechanism. We then focus specifically on the theoretical analysis of LQ games.
To formalize our guided policy optimization in LQ games, let $\check{K}$ be an arbitrary stabilizing feedback policy for the dynamics in (8). We consider the LQ games in Section III-B, but now with a cost for agent $i$, regulated by a deviation from the guiding control $\check{K}^{i}x_{t}$, given as
$\hat{J}_{\rho}^{i}$ $(K^{1},\dots,K^{N})$ (13)
${\scriptstyle:=}$
$\mathbb{E}_{x_{0}}\!\Big[\!\sum_{t=0}^{\infty}\big(x_{t}^{\top}Q^{i}x_{t}+{x}_{t}^{\top}{K}^{i\top}R^{i}{K}^{i}{x}_{t}\big)+\rho{x_{t}^{\top}(K^{i}-\check{K}^{i})^{\top}R^{i}(K^{i}-\check{K}^{i})x_{t}}\Big]$
${\scriptstyle=}$
${J}^{i}{(K^{1},\dots,K^{N})}+\rho\mathbb{E}_{x_{0}}\Big[\sum_{t=0}^{\infty}{x_{t}^{\top}(K^{i}-\check{K}^{i})^{\top}R^{i}(K^{i}-\check{K}^{i})x_{t}}\Big].$
We denote $\hat{K}^{*}=[\hat{K}^{1*},\dots,\hat{K}^{N*}]$ as the Nash equilibrium solution for this guiding-policy-regularized LQ game. We show that the introduced regularization term can resolve the instability of the MA-PG dynamics (7).
Proposition 1.
Under Assumption 1, let $K^{*}$ be a feedback Nash equilibrium policy, $\check{K}$ an arbitrary stabilizing feedback policy, and $\mathcal{K}$ the set of all stabilizing feedback policies. Then, we have
  1. 1.
    $w(K)$ is locally Lipschitz continuous on $\mathcal{K}$. There exists a radius $r$ such that $\forall K\in\mathcal{K}_{r}:=\{K\in\mathcal{K}:\|K-K^{*}\|_{2}^{2}\leq r\}$, the real parts of the eigenvalues of the Jacobian matrix $\nabla w(K)$ are lower-bounded by some $\lambda_{\min}$.
  2. 2.
    As we increase $\rho$, the regularized Nash equilibrium $\hat{K}^{*}$ will exhibit locally stable policy gradient dynamics, with $\lambda_{\min}>0$.
  3. 3.
    The bias of the regularized Nash equilibrium $\hat{K}^{i*}$ of each agent $i$ is characterized by
    $\Big\|$ $\hat{K}^{i*}-\left(\frac{\rho}{1+\rho}\check{K}^{i}+\frac{1}{1+\rho}K^{i*}\right)\Big\|_{2}$ (14)
    $\leq\frac{1}{1+\rho}\|{R^{i}}^{-1}\|_{2}\|B^{i\top}\|_{2}\big\|\hat{P}_{\hat{K}^{*}}^{i}\bar{A}_{\hat{K}^{*}}-P_{K^{*}}^{i}\bar{A}_{K^{*}}\big\|_{2},$
    where $\bar{A}_{\hat{K}}=A-\sum_{i=1}^{N}B_{i}\hat{K}_{i}$, and $\hat{P}_{\hat{K}}^{i}$ is given as the unique positive definite solution of $\hat{P}_{\hat{K}}^{i}={\bar{A}_{\hat{K}}}^{\top}\hat{P}_{\hat{K}}^{i}\bar{A}_{\hat{K}}+\hat{K}^{i}{}^{\top}R^{i}\hat{K}^{i}+Q^{i}+\rho{(\hat{K}^{i}-\check{K}^{i})}^{\top}R^{i}(\hat{K}^{i}-\check{K}^{i})$.
Proof.
We include the proof in Appendix. ∎
Figure 1: Policy guidance stabilizes infinite-horizon LQ game dynamics, enables convergence to a close neighborhood of the true Nash equilibrium—even if the guidance is wrong. The first two rows are the trajectories of the values of the two entries of $K^{1},K^{2}$ under policy gradient with different values of $\rho$ and a biased guiding policy $\check{K}$, which may come from an inaccurate dynamics model. A small $\rho$ ($\geq 0.001$) is sufficient to stabilize the policy gradient dynamics nearby the ground truth feedback Nash equilibrium $K^{*}$. However, as $\rho$ increases, the bias in the converged policies introduced by the guidance also grows.
Proposition 1 indicates that adding the guidance term from (13) with a sufficiently large $\rho$ stabilizes the policy gradient dynamics, guaranteeing the local exponential convergence of policy gradient methods to a stationary point even if the original policy gradient dynamics is unstable (i.e., $\lambda_{\min}<0$). Denote by $R\in\mathbb{R}^{m\times m}$ a diagonal block matrix where the $i$-th diagonal block is composed by $R^{i}$. This stability is primarily achieved because the term $(\rho\cdot R)$ shifts all eigenvalues of the pseudo-gradient’s Jacobian $(\nabla w(K)+\rho\cdot R)$ to the positive domain. However, the distance (bias) between the converged solution and the true Nash equilibrium depends jointly on the problem dynamics, $\rho$, and the quality of the guidance policy. We include a numerical study in Figure 1 to support this. Specifically, there are three cases:
  1. 1.
    When $\rho$ is small, we may not be able to stabilize the limit cycle behavior of the policy gradient dynamics if $(\nabla w(K)+\rho\cdot R)$ has an eigenvalue with a negative real part;
  2. 2.
    Increasing $\rho$ to an appropriately large value, we enhance the stability of the policy gradient dynamics, and the converged solution under the guided policy gradient dynamics could be close to the Nash equilibrium of the original game;
  3. 3.
    Further increasing $\rho$, we will enhance stability but the converged solution may deviate from the Nash equilibrium of the original game and biases more towards the guidance policy.
These observations suggest that $\rho$ trades off stability and bias. If the guidance matches the true Nash equilibrium, additional policy gradient updates become unnecessary; otherwise, the converged solution under the guided policy gradient dynamics can outperform an imperfect guidance policy $\check{K}$.
V Guiding Multi-Agent Policy Gradient in Nonlinear Games
In this section, we extend the idea of leveraging a model-based prior to stabilize policy gradients in LQ games to the context of nonlinear games. We introduce an efficient multi-agent guided policy search algorithm that incorporates this model-based prior for nonlinear game scenarios.
V-A Efficient computation of local guidance for MA-PG
Unlike the LQ scenario, where a stabilizing policy can be easily computed as guidance, it is generally hard to obtain guidance for MA-PG for nonlinear games due to the complexity of dynamics. Inspired by the iLQGames method [7], which argues that along the Nash equilibrium state and control trajectories of a nonlinear game, the Nash equilibrium of the locally constructed LQ game along those state and control trajectories can approximate the Nash equilibrium of the original nonlinear game well. Building upon this idea, we construct local LQ game approximations along the neural network control output and the resulted state trajectory, and then use the Nash equilibrium of those local LQ game approximations as reference signal to guide the learning of multi-agent policy gradients.
The key innovation of our approach is efficiently utilizing such a local guidance, rather than repeatedly computing iLQGames solutions until convergence at each sampled state. This guidance aggregates multi-agent gradient information, aligning policy updates across strategic agents toward a consistent Nash equilibrium and thereby stabilizing and accelerating MA-PG in settings with multiple equilibria arising from complex objectives.
To achieve this, we introduce the time-varying finite-horizon linear quadratic games, with a given initial state ${\bar{x}}_{0}$:
$\min_{\{x_{t}\}_{t=0}^{T+1},\{{u}_{t}^{i}\}_{t=0}^{T}}$ $\sum_{{t}=0}^{T}{x}_{t}^{\top}{Q}_{t}^{i}{x}_{t}+2{q}_{t}^{i\top}{x}_{t}+{u}_{t}^{i\top}{R}_{t}^{i}{u}_{t}^{i}+2{r}_{t}^{i\top}{u}_{t}^{i}$ (15)
s.t. ${x}_{{t}+1}={A}_{t}{x}_{t}+\sum_{i=1}^{N}{B}_{t}^{i}{u}_{t}^{i},\ {x}_{0}={\bar{x}}_{0},\ t\in[T]$
[…]
${A}_{t}=\nabla_{{\bar{x}}_{t}}{f}({\bar{x}}_{t},{\bar{u}}_{t}),$ ${B}_{t}^{i}=\nabla_{{\bar{u}}^{i}_{t}}{f}({\bar{x}}_{t},{\bar{u}}_{t}),$
${Q}_{t}^{i}=\frac{1}{2}\nabla^{2}_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $q_{t}^{i}=\frac{1}{2}\nabla_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$
${R}_{t}^{i}=\frac{1}{2}\nabla^{2}_{{\bar{u}}^{i}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $r_{t}^{i}=\frac{1}{2}\nabla_{{\bar{u}}_{t}^{i}}{c}^{i}({\bar{x}}_{x},{\bar{u}}_{t}).$
Rather than running iLQGames to convergence for every sampled state $\bar{x}_{0}$ during training, our method directly uses the first element of the Nash equilibrium control sequence $\{u_{t}^{*}\}_{t=0}^{T}$ from the local LQ approximation (15) as the guidance control for ${\bar{x}}_{0}$. Since the Nash equilibrium control $\{u_{t}^{*}\}_{t=0}^{T}$ of local LQ game (15) is defined in the local coordinate around the nominal control trajectory $\{{\bar{u}}_{t}\}_{{t}=0}^{T}$, the guidance control $\check{u}_{0}$ of ${\bar{x}}_{0}$, which is defined in the global coordinate, should be defined as $\check{u}_{0}:={\bar{u}}_{0}+u_{0}^{*}$. For agent $i\in[N]$, we define the mapping from $({\bar{x}}_{0},{\pi_{\theta}})$ to the guidance control $\check{u}_{0}^{i}$ as the guidance policy $\check{{\pi}}^{i}:({\bar{x}}_{0},{\pi_{\theta}})\to\check{u}^{i}_{0}$. With the model-based prior $\check{\pi}^{i}$, we introduce the guided policy stage cost of agent $i$:
$\hat{{c}}^{i}({x},u^{1},\dots,u^{N}):={c}^{i}({x},u^{1},\cdots,u^{N})+\rho\|u^{i}-\check{{\pi}}^{i}(x,{\pi_{\theta}})\|_{2}^{2}\vskip-3.00003pt$ (16)
where $\rho>0$ is the guidance weight.
Remark 1 (On the uniqueness of Nash equilibria).
When multiple Nash equilibria exist in local LQ games, enumerating all Nash equilibria is unnecessary and computationally excessive. Instead, we adopt the unique minimum L2-norm Nash equilibrium [8] as synchronized guidance across agents. Proposition 1 shows that such synchronized guidance can effectively stabilize MA-PG dynamics. Empirically, we find that this guidance enables our method to outperform prior MARL approaches and converge faster in these settings, as shown in Section VI.
Input: Neural network policies of each player $\{\pi_{\theta}^{i}\}_{i=1}^{N}$ and the neural network value functions $\{V_{\theta^{\prime}}^{i,\pi_{\theta}}\}_{i=1}^{N}$, learning rate $\eta$, list of guidance weights $\{\rho_{i}\}_{i=1}^{N_{e}}$, number of epochs $N_{e}$, number of gradient steps $N_{g}$, number of sampled rollout trajectories $M_{r}$, batch size $M_{b}$, rollout trajectory length $T_{r}$
Output: Optimized policies $\{\pi_{\theta}^{i}\}_{i=1}^{N}$
1
2 Set data buffer $D=\{\}$
3 for $k\leftarrow 1$ to $N_{e}$ do
4 for $j\leftarrow 1$ to $N_{g}$ do
5      Sample $M_{r}$ initial conditions $\{x_{0}^{(j)}\}_{j=1}^{M_{r}}$.
6      For each $j\in\{1,\dots,M_{r}\}$, simulate trajectories $\tau^{(j)}=\{x_{0}^{(j)}\!,u_{0}^{(j)}\!,\dots,x_{T_{r}+1}^{(j)}\}$, under the policy $\{{\pi_{\theta}}^{i}\}_{i=1}^{N}$.
7      Update data buffer, $D\leftarrow\{\tau^{(j)}\}_{j=1}^{M_{r}}$
8    for each agent $i\in[N]$ do
9         Sampling a batch of states $\{x^{(j)}\}_{j=1}^{M_{b}}$
10         Obtain the guidance $\{\check{u}^{(j)}\}_{j=1}^{M_{b}}$ for each $\{x^{(j)}\}_{j=1}^{M_{b}}$, as outlined in Section V-A
[…]
12        Optimizing value function via minimizing Bellman error [34]:
13       $\theta^{\prime}\leftarrow\theta^{\prime}-{\eta}\nabla\sum_{j=1}^{M_{b}}\|V_{\theta^{\prime}}^{i,\pi_{\theta}}(x^{(j)})-c^{i}(x^{(j)},\pi_{\theta}(x^{(j)}))-\gamma V_{\theta^{\prime}}^{i,\pi_{\theta}}(f(x^{(j)},\pi_{\theta}(x^{(j)})))\|_{2}$
14
15
Algorithm 1 Multi-agent guided policy search
V-B Multi-agent guided policy search for nonlinear non-cooperative dynamic games
Leveraging the policy guided cost function (16), we introduce our multi-agent guided policy search (MA-GPS) method in Algorithm 1. At each iteration of policy gradient, we first sample a number of initial states, simulate nominal trajectories and store them into the data buffer. Subsequently, for efficiently training, we sample a small batch of data from the data buffer, compute their guidance control, and then take one step policy gradient descent for each agent to improve the policy and value function estimations. We iterate until convergence. In practice, we could decay the guidance weight $\rho$ periodically. This allows MA-GPS to tolerate imperfect guidance and unleash the computational power of MA-PG to refine the sub-optimal policy.
VI Experiments
We evaluate our method against MA-PG methods in three examples: 1) a two-agent linear quadratic (LQ) game, 2) a platooning scenario involving three vehicles merging into a single lane, and 3) a 24-dimensional six-player basketball formation, where agents strategically move to desired positions based on their roles. Implementation details [^2] Code is available at https://github.com/qugechen/MA-GPS of experiments are included in Appendix.
In our deep RL experiment, we compare our method with IPPO [24], IPPO with L2 regularization (equivalent to zero-input guidance for pure variance reduction), MA-PPO [25] and MA-DDPG [29]. Our IPPO implementation is adapted for non-cooperative dynamic games, where each agent learns its own value function and policy based on its individual objective, unlike the cooperative setting in [24]. Concretely, each agent is equipped with a value and a policy network, both updated via simultaneous gradient play, as outlined in (5). IPPO and MA-PPO both employ entropy-driven exploration, making them competitive baselines, whereas MA-DDPG learns deterministic policies without it.
Figure 2: Three-vehicle platooning experiment. (a) Total reward during GPU training time. Note that the training time includes the computation time of the LQ game solutions used to guide the MA-GPS. MA-GPS achieves the highest reward and lowest variance compared to the other four methods. (b) Trajectories of the vehicles. The leader vehicle guides the other vehicles to the center lane. MA-GPS effectively enables merging into a lane, while IPPO’s Car 3 gets stuck within the same GPU computation time.
Figure 3: Six-player strategic basketball formation experiment. (a) Total reward during GPU training time. MA-GPS achieves a higher-performing policy more quickly than IPPO, MA-DDPG, and MA-PPO, with lower variance when measured by GPU training time. In contrast, naive L2 policy regularization reduces variance but slows IPPO’s convergence. These results suggest that the local LQ game approximations introduced in Section V-A can be computed efficiently, thereby accelerating MARL policy convergence. (b) Trajectories of Players. MA-GPS learns more effective policies than IPPO and MA-PPO within the same GPU computation time, as evidenced by the distinct formation of agents.
1) LQ game. As noted in [5], MA-PG methods in LQ games exhibit limit-cycle behaviors and fail to converge to Nash equilibria when the Jacobian of $w(K)$ in (9) has both positive and negative eigenvalues. We reproduce this phenomenon and study the stabilizing effect of our guidance regularization term. Using the system dynamics described in Appendix VIII-A, we solve for the ground truth Nash equilibrium policies $K^{1*},K^{2*}$ via the Lyapunov iteration from [5]. We introduce an artificial bias to the guiding policies $\check{K}^{1},\check{K}^{2}$ in the regularized cost (13). The resulting policy trajectories under varying regularization weights $\rho$ are shown in Figure 1.
Without guidance ($\rho=0$), policy gradients exhibit limit-cycle behavior and fail to converge. Introducing a small regularization weight $\rho$ effectively shifts all eigenvalues of $\nabla w(K)$ to positive real parts, stabilizing policy gradient dynamics, thus empirically validating Proposition 1. As illustrated in Figure 1, both agents’ policies converge to a neighborhood around their true Nash equilibria, despite imperfect guidance. Increasing $\rho$, however, enlarges the bias in converged policies relative to the ground truth. Hence, the optimal regularization weight $\rho$ is the smallest sufficient to stabilize convergence, minimizing the guiding policy bias.
2) Three-vehicle platooning. We evaluate MA-GPS in a nonlinear dynamic setting. Initially positioned in separate highway lanes, Car 1 guides Cars 2 and 3 to a target lane. Figure 2 compares MA-GPS with state-of-the-art MARL methods. For fairness, training progress is reported against GPU time. Since exact Nash equilibria are difficult to compute in nonlinear games, policy quality is assessed through total rewards and trajectory simulations. MA-GPS consistently accelerates convergence and reduces variance during training, achieving variance reduction comparable to L2 regularization but with minimal bias, and learning high-performance policies significantly faster than baselines.
3) Six-player strategic basketball formation.
We demonstrate that MA-GPS scales to larger numbers of agents and converges faster to high-performing policies than IPPO without guidance, measured in training time. To illustrate, we simulate a simplified basketball formation where both offense and defense teams consist of three agents: a shooting guard, a center, and a forward. Defensive agents aim to stay between their assigned opponent and the basket. On offense, the center positions near the basket ($r=1.5$), the shooting guard holds the three-point line ($r=4$) to prepare for shots, and the forward sets a screen by moving toward the defender covering the shooting guard.
The agents’ objectives include mild non-differentiabilities. For example, letting $p_{y}$ denote a player’s y-position, a term like $\max(p_{y},0)$ imposes a one-sided penalty with a kink at the baseline ($p_{y}=0$), discouraging players from crossing it. This illustrates that MA-GPS can handle objectives beyond perfectly smooth settings, broadening its applicability to a wider range of multi-agent tasks.
As shown in Figure 3, MA-GPS noticeably increases the convergence rate and reduces variance compared to other four methods. As observed in the trajectory plot, as the number of agents increases, it becomes increasingly difficult for all agents to explore and find an effective policy together.Our guidance provides an effective prior for all agents, helping them converge to a reasonable equilibrium.
Remark 2 (Comparing with finite-horizon iLQGames).
We measure wall-clock computation time $\tau$ for iLQGames with horizon length $T=20$ on two tasks: three-vehicle platooning ($\tau=0.0485\pm 0.0042\,\mathrm{s}$) and six-player basketball formation ($\tau=46.81\pm 4.18\,\mathrm{s}$). Because iLQGames uses finite-horizon formulations, its computational time scales quickly with horizon length, player count, and objective complexity. In contrast, MA-GPS shifts computation to offline training and directly solves the infinite-horizon case, producing real-time executable neural network policies ($\tau=0.0016\pm 0.0012\,\mathrm{s}$) for long-horizon strategic decision-making.
VII Conclusion
This paper introduced a multi-agent policy gradient method enhanced by model-based guidance, which stabilizes learning dynamics and accelerates convergence in non-cooperative dynamic games. The proposed approach mitigates the unstable limit-cycle behaviors that often arise in multi-agent policy gradient methods, thereby enabling the efficient learning of complex Nash equilibrium strategies.
Future work includes relaxing the assumption of an approximate model by learning dynamics models online and developing tighter theoretical bounds on the gap between guided and true Nash equilibria in LQ settings to better understand the impact of imperfect guidance on solution bias. Another direction is to unify MARL stabilization techniques, such as entropy-driven exploration, with model-based guidance to characterize how the balance between exploration and exploitation of game-theoretic structure affects the learning of Nash equilibrium policies.

### v3

@@
 In such cases, they become computationally expensive and may converge to inferior policies [9].
-These limitations highlight the complementary strengths of MARL and mode-based optimization, suggesting that a hybrid method could provide both scalability and stability.
+These limitations highlight the complementary strengths of MARL and model-based optimization, suggesting that a hybrid method could provide both scalability and stability.
 In this paper, we move beyond prior efforts that integrate such approaches for cooperative multi-agent control [10, 11] and ask: how can we integrate these two approaches to improve policy learning for non-cooperative dynamic games?
@@
 We propose a framework that incorporates this prior into the reward function to regularize MA-PG updates.
-In doing so, it offers a new mechanism for stabilizing MARL that outperforms existing techniques based on entropy-driven exploration [12, 13].
-Our method not only achieves faster convergence but also yields neural network policies that can be executed in real time for complex multi-agent dynamic games, where classical model-based optimization methods are computationally demanding for timely decision-making.
+This constitutes a new mechanism for stabilizing MARL, outperforming existing MARL stabilization approaches that primarily rely on entropy-driven exploration [12, 13].
+Beyond faster convergence, our method yields neural network policies that can be executed in real time for complex multi-agent dynamic games, where classical model-based optimization methods are often computationally demanding for timely decision-making.
 Our contributions are:
@@
 Model-based priors have been explored to reduce variance in RL [40].
-Unlike [40], which regularizes the policy output directly, our method incorporates the regularization into the reward function.
-Prior works such as [12, 13, 41] stabilize MA-PG in finite-horizon games using entropy-driven exploration.
+Unlike [40], which regularizes the policy output directly, our method incorporates regularization into the reward function.
+Moreover, prior works such as [12, 13, 41] stabilize MA-PG in finite-horizon games using entropy-driven exploration.
 In contrast, we show that suitable policy guidance ensures stability in infinite-horizon settings, which could be more practical for real-time decision-making since the optimal policy is time-invariant and avoids computing complex time-varying Nash equilibrium policies.
@@
 We can represent the MA-PG update as a dynamical system with respect to $\pi$}, often referred to as the MA-PG dynamics:
-${\pi}\leftarrow{\pi}-\eta\nabla w({\pi}).$ (7)
-III-B Non-cooperative LQ games
-In this subsection, we introduce a special class of non-cooperative games, LQ games [1], in order to present a relatively simple case where we can theoretically characterize the performance of policy gradient methods.
-In infinite-horizon LQ games, we consider
-$\min_{{\pi}^{i}}\$$\mathbb{E}_{x_{0}\sim{\mathcal{D}}}\Big[\sum_{{t}=0}^{\infty}{x}_{t}^{\top}{Q}^{i}{x}_{t}+{u}_{t}^{\top}{R}^{i}{u}_{t}^{i}\Big]$ (8)
-s.t. ${x}_{{t}+1}={A}{x}_{t}+{B}{u}_{t},\;{u}_{t}\sim{\pi}^{i}({x}_{t}),\;\forall t=0,1,\dots\vskip-2.5pt$
-where discount factor is set to 1. Note that for policies resulting in stable closed-loop dynamics, we do not need to introduce the discount. More formally, we consider the following assumption:
+${\pi}\leftarrow{\pi}-\eta w({\pi}).$ (7)
+III-B Non-cooperative linear quadratic games
+In this subsection, we introduce a special class of non-cooperative games known as linear–quadratic (LQ) games [1], a simplified setting that allows theoretical characterization of the performance of policy gradient methods.
+In infinite-horizon LQ games, we consider for each agent $i\in[N]$,
+$\min_{{\pi}^{i}}\$$\mathbb{E}_{x_{0}\sim{\mathcal{D}}}\Big[\sum_{{t}=0}^{\infty}{x}_{t}^{\top}{Q}^{i}{x}_{t}+{u}_{t}^{i\top}{R}^{i}{u}_{t}^{i}\Big]$ (8)
+s.t. ${x}_{{t}+1}={A}{x}_{t}+\sum_{i=1}^{N}{B}^{i}{u}_{t}^{i},\;{u}_{t}^{i}\sim{\pi}^{i}({x}_{t}),\;\forall t=0,1,\dots$
+where $A\in\mathbb{R}^{n\times n}$, $B^{i}\in\mathbb{R}^{n\times m_{i}}$, $Q^{i}\in\mathbb{R}^{n\times n}$ and $R^{i}\in\mathbb{R}^{m_{i}\times m_{i}}$, $\forall i\in[N]$.
+We consider the following assumption:
 Assumption 1.
@@
   2. 2.
-    As we increase $\rho$, the regularized Nash equilibrium $\hat{K}^{*}$ will exhibit locally stable policy gradient dynamics, with $\lambda_{\min}>0$.
+    As $\rho$ increases, the regularized Nash equilibrium $\hat{K}^{*}$ will exhibit locally stable policy gradient dynamics, with all eigenvalues of the regularized pseudo-gradient $(\nabla w(K)+\rho\cdot R)$ having positive real parts.
   3. 3.
@@
     $\leq\frac{1}{1+\rho}\|{R^{i}}^{-1}\|_{2}\|B^{i\top}\|_{2}\big\|\hat{P}_{\hat{K}^{*}}^{i}\bar{A}_{\hat{K}^{*}}-P_{K^{*}}^{i}\bar{A}_{K^{*}}\big\|_{2},$
-    where $\bar{A}_{\hat{K}}=A-\sum_{i=1}^{N}B_{i}\hat{K}_{i}$, and $\hat{P}_{\hat{K}}^{i}$ is given as the unique positive definite solution of $\hat{P}_{\hat{K}}^{i}={\bar{A}_{\hat{K}}}^{\top}\hat{P}_{\hat{K}}^{i}\bar{A}_{\hat{K}}+\hat{K}^{i}{}^{\top}R^{i}\hat{K}^{i}+Q^{i}+\rho{(\hat{K}^{i}-\check{K}^{i})}^{\top}R^{i}(\hat{K}^{i}-\check{K}^{i})$.
+    where $\bar{A}_{\hat{K}}=A-\sum_{i=1}^{N}B^{i}\hat{K}^{i}$, and $\hat{P}_{\hat{K}}^{i}$ is given as the unique positive definite solution of $\hat{P}_{\hat{K}}^{i}={\bar{A}_{\hat{K}}}^{\top}\hat{P}_{\hat{K}}^{i}\bar{A}_{\hat{K}}+\hat{K}^{i}{}^{\top}R^{i}\hat{K}^{i}+Q^{i}+\rho{(\hat{K}^{i}-\check{K}^{i})}^{\top}R^{i}(\hat{K}^{i}-\check{K}^{i})$.
 Proof.
@@
 Figure 1: Policy guidance stabilizes infinite-horizon LQ game dynamics, enables convergence to a close neighborhood of the true Nash equilibrium—even if the guidance is wrong.
-The first two rows are the trajectories of the values of the two entries of $K^{1},K^{2}$ under policy gradient with different values of $\rho$ and a biased guiding policy $\check{K}$, which may come from an inaccurate dynamics model.
-A small $\rho$ ($\geq 0.001$) is sufficient to stabilize the policy gradient dynamics nearby the ground truth feedback Nash equilibrium $K^{*}$.
+Each row corresponds to one of $K^{1}$ and $K^{2}$.
+In each plot, the two axes represent the entries $K^{i,1}$ and $K^{i,2}$ of the corresponding $K^{i}$, and the curves depict their trajectories under policy gradient with different values of $\rho$ and a biased guiding policy $\check{K}$ (potentially derived from an inaccurate dynamics model).
+A small $\rho$ ($\geq 0.001$) is sufficient to stabilize the policy gradient dynamics near the ground truth feedback Nash equilibrium $K^{*}$.
 However, as $\rho$ increases, the bias in the converged policies introduced by the guidance also grows.
@@
 Denote by $R\in\mathbb{R}^{m\times m}$ a diagonal block matrix where the $i$-th diagonal block is composed by $R^{i}$.
-This stability is primarily achieved because the term $(\rho\cdot R)$ shifts all eigenvalues of the pseudo-gradient’s Jacobian $(\nabla w(K)+\rho\cdot R)$ to the positive domain.
+This stability is achieved because the term $(\rho\cdot R)$ shifts all eigenvalues of the regularized pseudo-gradient’s Jacobian $(\nabla w(K)+\rho\cdot R)$ to the positive domain.
 However, the distance (bias) between the converged solution and the true Nash equilibrium depends jointly on the problem dynamics, $\rho$, and the quality of the guidance policy.
@@
   3. 3.
-    Further increasing $\rho$, we will enhance stability but the converged solution may deviate from the Nash equilibrium of the original game and biases more towards the guidance policy.
+    Further increasing $\rho$, we will enhance stability, but the converged solution may deviate from the Nash equilibrium of the original game and be biased more towards the guidance policy.
 These observations suggest that $\rho$ trades off stability and bias.
@@
 This guidance aggregates multi-agent gradient information, aligning policy updates across strategic agents toward a consistent Nash equilibrium and thereby stabilizing and accelerating MA-PG in settings with multiple equilibria arising from complex objectives.
-To achieve this, we introduce the time-varying finite-horizon linear quadratic games, with a given initial state ${\bar{x}}_{0}$:
+To achieve this, we introduce the time-varying finite-horizon linear quadratic games, with a given initial state ${\bar{x}}_{0}$ and $i\in[N]$:
 $\min_{\{x_{t}\}_{t=0}^{T+1},\{{u}_{t}^{i}\}_{t=0}^{T}}$ $\sum_{{t}=0}^{T}{x}_{t}^{\top}{Q}_{t}^{i}{x}_{t}+2{q}_{t}^{i\top}{x}_{t}+{u}_{t}^{i\top}{R}_{t}^{i}{u}_{t}^{i}+2{r}_{t}^{i\top}{u}_{t}^{i}$ (15)
-s.t. ${x}_{{t}+1}={A}_{t}{x}_{t}+\sum_{i=1}^{N}{B}_{t}^{i}{u}_{t}^{i},\ {x}_{0}={\bar{x}}_{0},\ t\in[T]$
-[…]
+s.t. ${x}_{{t}+1}={A}_{t}{x}_{t}+\sum_{i=1}^{N}{B}_{t}^{i}{u}_{t}^{i},\ {x}_{0}={\bar{x}}_{0},$
+$\forall t=0,\dots,T$
+The horizon length $T$ is user-defined and can be relatively short if computational resources are limited.
+In our experiments (Section VI), we find $T=10$ or $20$ sufficient for generating effective guidance, though the optimal choice depends on game-specific dynamics and costs.
+We propose training a neural network policy $\pi_{\theta}$} for nonlinear games via model-based guided policy gradients.
+Below, we describe the procedure for computing the reference trajectory used to obtain policy guidance.
+Consider a $T$-horizon nominal state trajectory $\{{\bar{x}}_{t}\}_{{t}=0}^{T+1}$ steered by the neural network control $\{{\bar{u}}_{t}\}_{{t}=0}^{T}$, with ${\bar{u}}_{t}\sim{\pi_{\theta}}({\bar{x}}_{t})$, ${\bar{x}}_{{t}+1}={f}({\bar{x}}_{t},{\bar{u}}_{t}^{1},\dots,{\bar{u}}_{t}^{N}),\ \ \forall t=0,\dots,T$.
+We linearize the dynamics and quadraticize the cost to derive a local LQ game approximation around the nominal trajectory $\{{\bar{x}}_{0},{\bar{u}}_{0},\cdots,{\bar{x}}_{T+1}\}$, for all $i\in[N]$,
 ${A}_{t}=\nabla_{{\bar{x}}_{t}}{f}({\bar{x}}_{t},{\bar{u}}_{t}),$ ${B}_{t}^{i}=\nabla_{{\bar{u}}^{i}_{t}}{f}({\bar{x}}_{t},{\bar{u}}_{t}),$
-${Q}_{t}^{i}=\frac{1}{2}\nabla^{2}_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $q_{t}^{i}=\frac{1}{2}\nabla_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$
-${R}_{t}^{i}=\frac{1}{2}\nabla^{2}_{{\bar{u}}^{i}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $r_{t}^{i}=\frac{1}{2}\nabla_{{\bar{u}}_{t}^{i}}{c}^{i}({\bar{x}}_{x},{\bar{u}}_{t}).$
+${Q}_{t}^{i}=\nabla^{2}_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $q_{t}^{i}=\nabla_{{\bar{x}}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$
+${R}_{t}^{i}=\nabla^{2}_{{\bar{u}}^{i}_{t}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}),$ $r_{t}^{i}=\nabla_{{\bar{u}}_{t}^{i}}{c}^{i}({\bar{x}}_{t},{\bar{u}}_{t}).$
 Rather than running iLQGames to convergence for every sampled state $\bar{x}_{0}$ during training, our method directly uses the first element of the Nash equilibrium control sequence $\{u_{t}^{*}\}_{t=0}^{T}$ from the local LQ approximation (15) as the guidance control for ${\bar{x}}_{0}$.
@@
 With the model-based prior $\check{\pi}^{i}$, we introduce the guided policy stage cost of agent $i$:
-$\hat{{c}}^{i}({x},u^{1},\dots,u^{N}):={c}^{i}({x},u^{1},\cdots,u^{N})+\rho\|u^{i}-\check{{\pi}}^{i}(x,{\pi_{\theta}})\|_{2}^{2}\vskip-3.00003pt$ (16)
+$\hat{{c}}_{\rho}^{i}({x},u^{1},\dots,u^{N}):={c}^{i}({x},u^{1},\cdots,u^{N})+\rho\|u^{i}-\check{{\pi}}^{i}(x,{\pi_{\theta}})\|_{2}^{2}\vskip-3.00003pt$ (16)
 where $\rho>0$ is the guidance weight.
@@
 5      Sample $M_{r}$ initial conditions $\{x_{0}^{(j)}\}_{j=1}^{M_{r}}$.
-6      For each $j\in\{1,\dots,M_{r}\}$, simulate trajectories $\tau^{(j)}=\{x_{0}^{(j)}\!,u_{0}^{(j)}\!,\dots,x_{T_{r}+1}^{(j)}\}$, under the policy $\{{\pi_{\theta}}^{i}\}_{i=1}^{N}$.
-7      Update data buffer, $D\leftarrow\{\tau^{(j)}\}_{j=1}^{M_{r}}$
+6      For each $j\in\{1,\dots,M_{r}\}$, simulate a trajectory $\xi^{(j)}=\{x_{0}^{(j)}\!,u_{0}^{(j)}\!,\dots,x_{T_{r}+1}^{(j)}\}$, under the policies $\{\pi_{\theta}^{i}\}_{i=1}^{N}$.
+7      Update data buffer, $D\leftarrow\{\xi^{(j)}\}_{j=1}^{M_{r}}$
 8    for each agent $i\in[N]$ do
 9         Sampling a batch of states $\{x^{(j)}\}_{j=1}^{M_{b}}$
-10         Obtain the guidance $\{\check{u}^{(j)}\}_{j=1}^{M_{b}}$ for each $\{x^{(j)}\}_{j=1}^{M_{b}}$, as outlined in Section V-A
-[…]
-12        Optimizing value function via minimizing Bellman error [34]:
-13       $\theta^{\prime}\leftarrow\theta^{\prime}-{\eta}\nabla\sum_{j=1}^{M_{b}}\|V_{\theta^{\prime}}^{i,\pi_{\theta}}(x^{(j)})-c^{i}(x^{(j)},\pi_{\theta}(x^{(j)}))-\gamma V_{\theta^{\prime}}^{i,\pi_{\theta}}(f(x^{(j)},\pi_{\theta}(x^{(j)})))\|_{2}$
-14
-15
+10         Optimizing policy: $\theta\leftarrow\theta-{\eta}\nabla_{\theta}\sum_{j=1}^{M_{b}}\Big[\hat{{c}}_{\rho_{k}}^{i}\big(x^{(j)},\pi_{\theta}(x^{(j)})\big)+\gamma V_{\theta^{\prime}}^{i,\pi_{\theta}}\Big(f\big(x^{(j)},\pi_{\theta}(x^{(j)})\big)\Big)\Big]$
+11        Optimizing value function via minimizing Bellman error [34]: $\theta^{\prime}\leftarrow\theta^{\prime}-{\eta}\nabla_{\theta^{\prime}}\sum_{j=1}^{M_{b}}\Big\|V_{\theta^{\prime}}^{i,\pi_{\theta}}(x^{(j)})-\hat{{c}}_{\rho_{k}}^{i}\big(x^{(j)},\pi_{\theta}(x^{(j)})\big)-\gamma V_{\theta^{\prime}}^{i,\pi_{\theta}}\Big(f\big(x^{(j)},\pi_{\theta}(x^{(j)})\big)\Big)\Big\|_{2}$
+12
+13
 Algorithm 1 Multi-agent guided policy search
@@
 As noted in [5], MA-PG methods in LQ games exhibit limit-cycle behaviors and fail to converge to Nash equilibria when the Jacobian of $w(K)$ in (9) has both positive and negative eigenvalues.
-We reproduce this phenomenon and study the stabilizing effect of our guidance regularization term.
-Using the system dynamics described in Appendix VIII-A, we solve for the ground truth Nash equilibrium policies $K^{1*},K^{2*}$ via the Lyapunov iteration from [5].
+We reproduce this phenomenon in an LQ game example, detailed in the Appendix, and analyze the stabilizing effect of the proposed guidance regularization term.
+We solve for the ground truth Nash equilibrium policies $K^{1*},K^{2*}$ via the Lyapunov iteration from [5].
 We introduce an artificial bias to the guiding policies $\check{K}^{1},\check{K}^{2}$ in the regularized cost (13).
@@
 On offense, the center positions near the basket ($r=1.5$), the shooting guard holds the three-point line ($r=4$) to prepare for shots, and the forward sets a screen by moving toward the defender covering the shooting guard.
-The agents’ objectives include mild non-differentiabilities.
-For example, letting $p_{y}$ denote a player’s y-position, a term like $\max(p_{y},0)$ imposes a one-sided penalty with a kink at the baseline ($p_{y}=0$), discouraging players from crossing it.
-This illustrates that MA-GPS can handle objectives beyond perfectly smooth settings, broadening its applicability to a wider range of multi-agent tasks.
+We note that some agents’ objectives include mild non-differentiabilities.
+For instance, a player’s objective may contain a term like $\max(p_{y},0)$, where $p_{y}$ denotes that player’s y-position.
+This one-sided penalty introduces a kink at the baseline ($p_{y}=0$), discouraging the player from crossing it.
+Our experiments indicate that MA-GPS can accommodate such objectives beyond perfectly smooth settings, thereby broadening its applicability to a wider range of multi-agent tasks.
 As shown in Figure 3, MA-GPS noticeably increases the convergence rate and reduces variance compared to other four methods.
-As observed in the trajectory plot, as the number of agents increases, it becomes increasingly difficult for all agents to explore and find an effective policy together.Our guidance provides an effective prior for all agents, helping them converge to a reasonable equilibrium.
+As observed in the trajectory plot, as the number of agents increases, it becomes increasingly difficult for all agents to explore and find an effective policy together.
+Our guidance provides an effective prior for all agents, helping them converge to a reasonable equilibrium.
 Remark 2 (Comparing with finite-horizon iLQGames).
-We measure wall-clock computation time $\tau$ for iLQGames with horizon length $T=20$ on two tasks: three-vehicle platooning ($\tau=0.0485\pm 0.0042\,\mathrm{s}$) and six-player basketball formation ($\tau=46.81\pm 4.18\,\mathrm{s}$).
+We measure wall-clock computation time $\tau$ for iLQGames with horizon length $T$ on two tasks: three-vehicle platooning ($T=20,\tau=0.0485\pm 0.0042\,\mathrm{s}$) and six-player basketball formation ($T=100,\tau=0.64\pm 0.15\,\mathrm{s}$).
 Because iLQGames uses finite-horizon formulations, its computational time scales quickly with horizon length, player count, and objective complexity.

### v4

@@
 Model-based solutions to non-cooperative games.
-Computing Nash equilibria of non-cooperative dynamic games is NP-hard [14]; however, under smoothness assumptions on dynamics and costs, they can be solved in polynomial time [1].
+Computing Nash equilibria of non-cooperative games is computationally challenging [7]; however, under smoothness assumptions on dynamics and costs, they can be solved in polynomial time [3].
 For instance, LQ games enable efficient computation, with feedback Nash equilibria obtained via coupled Riccati equations or KKT conditions [20].
@@
 In contrast, learning game-theoretic equilibria [24] (e.g., Nash equilibria) with deep RL remains challenging due to coupled dynamics and conflicting objectives.
-Policy gradient methods converge reliably only in special classes such as two-agent zero-sum games [30, 31] and potential games [32], where agents’ costs are symmetric.
+Policy gradient methods and their variants converge reliably in two-agent zero-sum games [42, 6, 34] and potential games [15], where agents’ costs are symmetric.
 In general non-cooperative games, however, they may converge to limit cycles rather than a Nash equilibrium [26].
@@
 We denote $\hat{K}^{*}=[\hat{K}^{1*},\dots,\hat{K}^{N*}]$ as the Nash equilibrium solution for this guiding-policy-regularized LQ game.
-We show that the introduced regularization term can resolve the instability of the MA-PG dynamics (7).
+We show that the introduced regularization term can resolve the instability of the MA-PG dynamics (7), thereby guaranteeing local exponential convergence to a stationary point of the regularized game.
 Proposition 1.
@@
 These observations suggest that $\rho$ trades off stability and bias.
-If the guidance matches the true Nash equilibrium, additional policy gradient updates become unnecessary; otherwise, the converged solution under the guided policy gradient dynamics can outperform an imperfect guidance policy $\check{K}$.
+Moreover, if the guidance matches the true Nash equilibrium, additional policy gradient updates become unnecessary; otherwise, the converged solution under the guided policy gradient dynamics can outperform an imperfect guidance policy $\check{K}$.
 V Guiding Multi-Agent Policy Gradient in Nonlinear Games
@@
 Inspired by the iLQGames method [12], which argues that along the Nash equilibrium state and control trajectories of a nonlinear game, the Nash equilibrium of the locally constructed LQ game along those state and control trajectories can approximate the Nash equilibrium of the original nonlinear game well.
-Building upon this idea, we construct local LQ game approximations along the neural network control output and the resulted state trajectory, and then use the Nash equilibrium of those local LQ game approximations as reference signal to guide the learning of multi-agent policy gradients.
+Building upon this idea, we construct local LQ game approximations along the neural network control output and the resulting state trajectory, and then use the Nash equilibrium of those local LQ game approximations as reference signal to guide the learning of multi-agent policy gradients.
 The key innovation of our approach is efficiently utilizing such a local guidance, rather than repeatedly computing iLQGames solutions until convergence at each sampled state.
@@
 We evaluate our method against MA-PG methods in three examples: 1) a two-agent linear quadratic (LQ) game, 2) a platooning scenario involving three vehicles merging into a single lane, and 3) a 24-dimensional six-player basketball formation, where agents strategically move to desired positions based on their roles.
+Although the last two examples fall outside Assumption 1, our method remains effective, suggesting that the stability analysis of Theorem 1 may extend beyond the LQ setting.
 Implementation details [^2] Code is available at https://github.com/qugechen/MA-GPS of experiments are included in Appendix.
@@
 Our experiments indicate that MA-GPS can accommodate such objectives beyond perfectly smooth settings, thereby broadening its applicability to a wider range of multi-agent tasks.
-As shown in Figure 3, MA-GPS noticeably increases the convergence rate and reduces variance compared to other four methods.
+As shown in Figure 3, MA-GPS noticeably increases the convergence rate and reduces variance compared to other four methods, particularly during the early stages of training, highlighting the stabilizing effect of the proposed model-based guidance.
 As observed in the trajectory plot, as the number of agents increases, it becomes increasingly difficult for all agents to explore and find an effective policy together.